OpenStudy (anonymous):

f 3a + 2b = 7 and 2a + 2b = 9 , what is the value of ?

OpenStudy (aripotta):

the value of what?

OpenStudy (anonymous):

5a + 4b over 2

OpenStudy (anonymous):

I got 8 as my answer

OpenStudy (anonymous):

I just wanted to know if i was correct

OpenStudy (anonymous):

A = -2 B= 13/2 3(-2) + 2(13/2) = 7 and 2(-2) + 2(13/2) = 9

OpenStudy (anonymous):

oh so i was close lol

OpenStudy (anonymous):

thanks!